2x + 4y = -2
3x - 2y = -5

Answer by uma(370)

2x + 4y = -2 ---------------(1)
3x - 2y = -5 ------------------(2)
(2) multiplied by 2 gives, 6x - 4y = -10 --------(3)
2x + 4y = -2 ------------(1)

Adding the above equations,
8x = - 12
==> x = -12/8
= -3/2
Plugging this in (1), 2(-3/2) + 4y = -2
-3 + 4y = -2
4y = -2 + 3
4y = 1
y = 1/4
Thus x = -3/2 and y = 1/4

Good Luck!!!